2.56grams of sulphur  in 100ml solution  shows osmotic  pressure  of 2.463atm at 300k.how many sulphur  atoms are associated in colloidalsol? 

Grade:12

1 Answers

Vikas TU
5 years ago
We need to calculate first the observed molar mass.
O.P = W*R*T/(M*V)
M = WRT/((O.P)*V)
Put the values,
M = (2.56)*0.0821*300/(2.463)*(0.1)
    = 256 grams/mol.
 
Number of atomes = > 256/32 => 8.